OpenStudy (anonymous):

PLEASE HELP!!!! I WILL MEDAL AND FAN!!!! What is the slope of the graph of the equation below? x+y=5 A -1 B 1 C -5 D 5

OpenStudy (anonymous):

There isn't a graph to go with this

jimthompson5910 (jim_thompson5910):

The idea is to solve for y x+y = 5 x+y-x = 5-x ... Subtract x from both sides. y+x-x = -x+5 y+0x = -x+5 y+0 = -x+5 y = -x+5 Your steps don't need to be this detailed. I put extra detail in there to be more formal

jimthompson5910 (jim_thompson5910):

The equation `y = -x+5` is the same as `y = -1x+5`

jimthompson5910 (jim_thompson5910):

The equation `y = -1x+5` is in the form `y = mx+b`